if ((a)^(2))=((b)^(3))=((c)^(5))=((d)^(6)) , then log abc , base d

Grade:12

1 Answers

Prasanjeet Mohanty Mohanty
14 years ago

Let a^2=b^3=c^5=d^6=k. Then a=k^1/2   b=k^1/3  c=k^1/5  d=k^1/6. 

then abc=k^31/30. using the properties of logaritms we can solve it to find the answer.on solving it comes to be 31/5.